the domain of defination of f(x) =log 2 (x+3)/x 2 +3x+2


  • the domain of defination of f(x) =log2(x+3)/x2+3x+2


 

Grade:12

1 Answers

Khushboo Singh
13 years ago

f(x)=log2(x+3)/(x2+3x+2)  =log2(x+3)/(x+1)(x+2)

x+3>0  =>x>-3   -----(1)

(x+1)(x+2) ≠ 0

x ≠-1,-2  ----(2)

from eq (1) and (2)

domain is (-3,∞)/{-1.-2}